RaNdOmUsEr3311:

Hi users of question cove. I have a discussion thing for math that I need help with. Here is the prompt: What do you consider to be the advantages of the numerical values? What do you consider to be the advantages of the visual representations? Which do you prefer working with, and why? I just need someone to answer what the advantages of the numerical values and the advantages of the visual representations are. The last part of the prompt I can do.

Vocaloid:

Please do not copy-paste my response. This is just to help you think of ideas about this topic. Let’s use multiplication as an example You could use numerical values to write 4*5 = 20, or use visuals to draw 4 groups of 5 or 5 groups of 4 objects and show that there are 20 total objects. I would argue that the visual is good for learning/showing basic concepts, while the numeric system is more compact and easier to do calculations with. Imagine instead multiplying 40 * 50. Using the numeric system we can simply see that 40 * 50 = 4 * 10 * 5 * 10 or 4 * 5 * 100 so it’s simply 100 times our previous result. However, having to draw this out would be more complicated, and the more complicated the numbers/arithmetic is, the harder it would be to visualize/draw it.

RaNdOmUsEr3311:

thank you so much!